OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1

OpenStudy (anonymous):

(3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (compassionate):

There are two methods for solving a system of linear equations. Elimination and Substitution. Which one do you think we should perform?

OpenStudy (anonymous):

y =1 + 2x put this equation in 2nd eq. and solve for x then use that value of x in first eq. to get the value of y

OpenStudy (anonymous):

what

OpenStudy (compassionate):

-2x + y = 1 -4x + y = -1 Okay, so we're going to solve by elimination. First, what do we need to multiply either of the two equations by in order to get the x terms to cancel each other?

OpenStudy (anonymous):

um

OpenStudy (compassionate):

-2x + y = 1 -4x + y = -1 If we multiply the first equation by -2, it will cancel. As shown -2(-2x + y = 1) -4x + y = -1 4x + y = 1 -4x + y = -1 Now you can add both the equations together, and the x terms should cancel. Then you can solve for y. Once you have y, plug it into either equation and solve for x. This will give you your final solution.

OpenStudy (anonymous):

y=1 y=-1 thats after the x is canceled right but i dont know what i do with that

OpenStudy (compassionate):

Simply add both of them together 4x + -4x + y + y = 1 + -1 Did you see what I did? Just simply adding both equations together. You'll notice the x terms will cancel. And we have 2y = 0 Now divide 2 by 0 .

OpenStudy (anonymous):

0

OpenStudy (compassionate):

Right. So when y = 0, we are done. There are No Solutions. When 0 = 0, there are infinite solutions. The more you do System of Equations, the more you'll get it. Want to do some practice problems? Do you have any similar problems?

OpenStudy (anonymous):

i have similar problems and thanks im kinda getting the hang of it but i think i should do more of them too

OpenStudy (compassionate):

Okay. Go ahead and post and I'll help you here. No need to make a new question. I don't care about medals or anything.

OpenStudy (anonymous):

wait for the last question none of the choices were no solutions

OpenStudy (compassionate):

What are they.

OpenStudy (anonymous):

(3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

(3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (compassionate):

-2x + y = 1 -4x + y = -1 -2(-2x + y = 1) -4x + y = -1 4x -2y = -2 -4x + y = - 1 Now we add them together 4x - 4x + y - 2y = -2 - 1 y - 2y = -2 - 1 -y = -3 y = 3 Now take 3 and plug it into either equation to solve for x -2x + y = 1 -2x + 3 = 1 -2x = -2 x = 1 So we have (1, 3) Whoops! Hehe. I see what I did earlier. Do you understand what I just did? I basically solved for y, then plugged y in and solved for x.

OpenStudy (anonymous):

ohh that makes sense thanks

OpenStudy (compassionate):

Have any more? The more we do the better you'll get.

OpenStudy (anonymous):

i have like 4 more but i think after like 1 or 2 i will probably understand

OpenStudy (compassionate):

Lets do another one.